### v2

@@ Abstract
keywords:
Deep learning, Dark-field, Near-field, Far-field, Single plasmonic nanocavity, Consistency, Predicting spectra from images, Classification
@@ Introduction
Plasmonic nanostructures, particularly those forming nanogaps, support localized surface plasmons that confine light to deep-subwavelength volumes comparable to molecular and atomic dimensions. [1, 2, 3, 4, 5]. By integrating low-dimensional materials into these nanogaps, they enable probing of electronic (e.g., photoluminescence) [6], vibrational (e.g., surface-enhanced Raman scattering) [7, 4], and other linear and nonlinear optical properties of materials [8, 9], with atom-level sensitivity [10], introducing a fascinating field as ‘extreme nanophotonics[3, 11]. This high sensitivity arises from the extremely enhanced nanoscopic light-matter interactions supported by nanostructure’s quasinormal modes [12], making resonance–material spectral matching a fundamental prerequisite for nanophotonic applications.
@@
However, this extreme localization also amplifies contributions from local nanomorphological features—such as nonuniformity, roughness, defects, surface atomic-level dynamics, and pico-cavities—inevitably introducing ‘noise’ and compromising reproducibility [13, 14, 15, 16, 12, 17, 4, 18, 19]. This intrinsic variability [20, 3] poses a paradox—enhancing fields for exceptional single-particle performance can inherently undermine spectral consistency and reproducibility across ensembles, constituting a crucial bottleneck in extreme nanophotonics. One general way of circumventing this is to identify and pre-screen the target (e.g., spectrally-uniform, with desired resonances) nanoparticles from countless particles ($\simeq 10^{10}$ per milliliter in the solution) drop-casted on the substrate during the sample preparation, and conduct experiments only on that subset. Conventionally, this screening process has relied on manual expertise combined with imaging and spectroscopic measurements, rendering it inefficient and labor-intensive.
@@
Confocal dark-field (DF) microscopy has long been a cornerstone for single-nanoparticle characterization. Resonances and sample uniformity are often inferred empirically from their images at high throughput, assuming that the color distribution of Airy patterns directly reflects the energy and lineshape of the plasmonic modes. However, human vision cannot reliably resolve subtle chromatic variations, and the information-compressed three-digit RGB output of commercial cameras (400–700 nm) lacks sufficient spectral detail and fails to capture lower-energy resonances. Multiple resonances with distinct far-field patterns further complicate the problem [21, 22]. More precise techniques, such as DF mapping, acquire spatially resolved spectra via point-scanning, representing a form of hyperspectral imaging (HSI) [23, 24, 25, 26, 27, 28]. Such sequential acquisition is unsuitable for large-field detection, which can be mitigated with alternative HSI strategies, including line-scanning (push-broom), wavelength-scanning [23], and snapshot techniques [28]. Recent advances, including compressive sensing [29, 30, 31, 32], have further improved acquisition speed. However, HSI methods inherently face a resolution–throughput trade-off: higher spectral resolution requires narrower channels, reducing photon flux per channel [33, 27]. This lowers the signal-to-noise ratio and, unless compensated by stronger illumination or more sensitive detectors, demands longer exposure times—especially critical for weak signals such as single nanoplasmonic structures. They also entail costly and specialized hardware, large data storage and copious processing requirements, and rigorous calibration protocols [24, 34]. In contrast, confocal DF microscopy combined with RGB imaging remains economical and effective, while readily compatible with emerging techniques (e.g., machine learning, compressive sensing). However, to fully leverage this approach for high-throughput, precise acquisition of plasmonic nanostructures, the complex and nonlinear correlations between DF images and spectra—particularly beyond the RGB camera’s physical limits—must be revealed.
@@
In this regard, deep learning (DL) has emerged as a transformative tool for decoding complex photonic interactions [35]. It has demonstrated the ability to learn intricate light–matter correlations through inverse design of plasmonic devices and metasurfaces [36, 37, 38, 39, 40, 41, 42, 43]. When integrated with microscopic and nanoscopic techniques, DL has also shown success in noise suppression, signal processing, nanoparticle identification, and super-resolution imaging, which resolves variations and uncertainties at the subwavelength scale [44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Overall, DL is particularly valuable for managing uncertainties in experimental and fabrication processes—such as material inconsistencies, structural imperfections, and environmental fluctuations.
@@
Furthermore, by leveraging heteroscedastic loss, the SPARX model could estimate the uncertainty in spectral predictions for individual nanoparticles, effectively quantifying the prediction error. This correlation allows for further refinement in the selection process. Moreover, as a powerful complementary screening approach, we develop a SPARX classifier capable of directly identifying and differentiating nanoparticle shapes from DF RGB images—showcasing an inference capability beyond the reach of conventional spectral acquisition methods. All in all, our approach achieves a $10^{2}$$10^{4}$-fold acceleration in characterization speed compared with spectral acquisition, enabling real-time classification of nanostructures in the future. By replacing spectrometer-dependent workflows with camera-based deep learning, we establish a scalable platform for next-generation optical characterization, nanophotonic device engineering, and biosensor development, where both spectral precision and high-throughput characterization have proven indispensable.
Results
@@
Among all plasmonic systems, the nanoparticle-on-mirror (NPoM) configuration (Figure 1(a)) has been widely observed as one of the most versatile platforms, offering easily controllable nanogaps for enhancing optical phenomena [3]. Therefore, we select NPoM as a representative platform for demonstration, measuring 12,000 gold NPoMs to compile a dataset that captures the underlying complexity. In our experiment, the NPoM system was formed by drop-casting 80 nm cetyltrimethylammonium (CTAC)-capped gold nanoparticles onto a gold mirror (see Methods for sample preparation). The nanogap was defined by a 1-2 nm thick CTAC molecule layer. From electromagnetic theory, it is well established that the size and shape of the nanoparticles [15, 18], small variations in the roughness of the metallic surface [14, 13], and the morphology of the gap between the nanoparticles and the metallic substrate (e.g., facets size and adatoms) [19, 12] can drastically alter the resonance behavior of the system.
To visualize these effects, we performed finite element method (FEM) simulations for the DF scattering spectra of the NPoMs using the commercial software COMSOL Multiphysics (see Methods for details). To avoid exhaustive geometric cases discussed above, our simulations focus on two primary geometric factors varying at the nanometer scale: the facet size, and the nanogap thickness between the nanoparticle and the substrate. The results, presented in Figure 1(b), clearly show that even minor modifications to these geometrical parameters lead to significant shifts in the resonances. Increasing the facet size may redshift the resonances, which can be related to the polyhedral morphology of the metal’s crystalline structure and light-induced atomic migration [15]. The nanogap thickness is another crucial factor that may substantially shift the resonances, since the nanogap plasmons have been proven to be able to resolve sub-picometer level of thickness variation [54]. In our test sample, the gap thickness can inherently vary on the level of a few angstroms.
@@
The attempt to uncover the correlations among different orders of the plasmonic resonances can be dated back to the Mie theory [55], where an analytical solution of the scattering from a spherical nanoparticle can be derived by solving the Maxwell equations. Each Mie resonance (i.e., dipole, quadrupole, etc.) can be intercorrelated through the spherical harmonics. When coupled with its own image through a mirror, the NPoM plasmonic nanogap can give rise to a richer variety of the hybridized modes [56], with the resonance of $mn$-order mode $\lambda_{mn}$ given by the cylindrical Bessel functions [19]: $\lambda_{mn}=\frac{\pi wn_{\mathrm{eff}}}{J_{mn}-\phi}$, where $J_{mn}$ is $n$-th root of the $m$-th order Bessel function. $n_{\rm eff}$ is the effective refractive index of the metal-insulator-metal junction, and $\phi$ is a proper reflection phase. However, this simplified model can only predict the cavity-like resonances of a perfectly spherical nanoparticle with a single ideal circular facet at the bottom gap, with uncertainties such as polyhedral shapes, roughness, nonlocal and quantum effects, and others mentioned above falling outside its scope. Thus, using a simplified, classical analytical solution to extend the spectra of a realistic system beyond the detection limit can be challenging.
In contrast, given sufficiently representative datasets, DL methods excel at resolving complex and nonlinear correlations with high degrees of freedom. In fact, our SPARX model has managed to decode the broadband spectrum varying from 500-1000 nm from the images and therefore could select the spectra with the minimal predicted MAE compared with the reference. It hit 3 out of 10 actual best choices as shown in Figure 1 (c), breaking away from the color-based selection rule. Nonetheless, the rest 7 “missed” selections also have a great match with the reference peak, showing a nearly overlapped mean value (Figure 1 (d)) with low standard deviation. Additionally, Figure 1(e) summarizes the mean, standard deviation, minimum, and maximum values of the resonance peak location differences with respect to the reference particle. Again, this analysis shows that the SPARX’s selection closely aligns with the actual best selection, exhibiting similar shifts and variances. This reinforces the capability of deep learning in reliably identifying and predicting nanoparticles with optimal spectral properties, outperforming human intuition. Next, we will explain the process of preparing the datasets and training the SPARX model.
@@
We perform principal component analysis (PCA) on both the DF images and the processed spectra. PCA reduces high-dimensional data by projecting it onto new orthogonal axes—principal components—that capture most of the variance (i.e., the directions in which the data varies the most or has the most spread). See Methods for data processing and multivariate analysis for more details. The variance explained by the first 128 PCA components for both datasets is shown in Figure 2(b). Figures 2(c) and (d) display the first four PCA components of the DF images and spectral data, respectively. In Figure 2 (c), each PCA component is presented with its corresponding red, green, and blue channels. A notable observation is the correlation between the size of extracted features and the diffraction limit of each color channel. For instance, in PC1 of the DF image (Figure 2(c), top row), the dark central region in RGB combined plot (where the data is near zero) appears in different colors across the RGB image, but its size decreases from red to blue, consistent with diffraction-limited resolution. Similar trends are observed in other components, such as the concentric ring structures in PC2, which also exhibit size reduction across color channels. On the other hand, the patterns of the PC1–4 for the DF images clearly present the spherical harmonics (e.g., $l=0$ for PC1, $l=1$ for PC2–4), inferring the NPoM’s point scattering nature in the Airy pattern [21]. These observations demonstrate that our analysis based on PCA captures meaningful physical features, revealing both the optical and structural characteristics of the system. Furthermore, using the PCA of DF images we are able to access scattering information and potentially spectral signatures.
@@
From the spectral PCA components in Figure 2(d), PC1—which captures the largest amount of variance—primarily reflects peak shift behavior. It shows higher values when the spectral peak shifts toward the red (1000 nm) or blue (400 nm) regions, with a local minimum around 800 nm. This behavior aligns with the simulations in Figure 2(b), where minor geometric variations, such as changes in facet size or nanogap thickness, lead to significant spectral shifts. The remaining PCs capture finer spectral details, with alternating positive peaks and negative dips across various wavelengths. Thus, our analysis based on the PCA of DF spectra not only reveals spectral variations but also encodes geometric information.
@@
To explore the correlation between DF images and spectra, we project both datasets into lower-dimensional PCA spaces and visualize the results in Figure 2(e). The density distribution of the spectral data along PC1 reveals multiple subpopulations, suggesting that this component captures distinct spectral characteristics. In the 2D PCA projection of DF images, two well-separated clusters emerge (middle panel of Figure 2 (e)). When coloring the DF image projections using different criteria (see Methods: Multivariate Analysis), clear patterns can emerge. First, using spectral PC1 projection reveals an alignment between DF image features and spectral variation. Second, using sample indices—reflecting the acquisition timeline—demonstrates the consistency of measurements over time. Third, spectral features, namely the peak location (wavelength $\lambda$) and total spectral energy, further validate that the image-based clustering corresponds to meaningful spectral characteristics. These patterns highlight the physical relevance of the clusters found in our analysis (Figure 2(e))
@@
To further support these observations, we applied Uniform Manifold Approximation and Projection (UMAP), a nonlinear dimensionality reduction technique, to both spectral and DF image data, as shown in SI Figure S2. The spectral data is embedded into a one-dimensional UMAP space, while the DF images are projected into two dimensions. By coloring the DF UMAP projection using various spectral-derived features—including UMAP values of the spectra, peak positions, and total spectral energy—we uncover strong visual correlations between the DF and spectral modalities (2D maps in SI Figure S2). An interesting outcome is the appearance of multiple distinct clusters in the DF UMAP space (in contrast to PCA projection in Figure 2 (e)), which exhibit clear correspondence to specific spectral features. For example, the cluster located at the bottom right of the projection space shows a color gradient where resonance wavelength and total spectral energy vary in opposite directions—i.e., as the resonance red-shifts, the intensity decreases. Additionally, when coloring the projections based on sample index—correlated with the data acquisition time—we observe no discernible pattern, confirming the high reproducibility and stability of the measurements across multiple days. This improved clustering allows us to reveal the rich spectral information that DF images inherently carry, which is typically inaccessible in traditional DF imaging analysis.
@@
In contrast, the intensity distribution, shown in the right panel of Figure 3(b), presents a distinctly different behavior. The overall population density plot of resonance intensity (shown in black) displays three distinct peaks, indicating the existence of three subpopulations. Interestingly, the core group is primarily composed of the subpopulation with the lowest resonance intensities. The remaining two subpopulations with higher intensities are predominantly classified as outliers. Interestingly, the spectra with modest intensity ($\simeq$2.6 arb. units), which contain the largest number of samples, are most recognized as the core data that SPARX is most confident in. This may be attributed to their generally more uniform and well-behaved geometries, with reduced randomness stemming from previously discussed factors. In contrast, the higher-intensity subpopulations are treated as outliers by SPARX, likely due to more complex features such as sharp polyhedral facets, nonuniform gaps, or even picocavities. These features have more randomness in general, and are more difficult to characterize using only DF images, leading to increased uncertainty in their prediction. In fact, this also aligns with the empirical intuition in plasmonics and surface-enhanced Raman scattering: extreme enhancement with high intensity often compromises robustness and consistency. Nanoparticles exhibiting such exceptionally high intensities remain rare outliers relative to the overall population. Remarkably, our SPARX model captures this trade-off by analyzing the associated uncertainty. It is important to note that while our heteroscedastic model is capable of identifying such complex cases from DF images, accurately reconstructing their spectral details remains challenging due to the limitations of the input modality. Detection of these outliers should not be conflated with precise spectral prediction.
@@
Beyond NPoM, we further validated the generality of SPARX on a more complex plasmonic system—nanopatch antennas, composed of single $\sim$75 nm gold nanocubes on a mirror (i.e., NCoM, see SI Section S5, Figs. S6, S7). Unlike nanospheres, nanocubes produce more diverse spectral responses, as geometric variations can generate multiple resonances associated with different gap sizes and bottom facet geometries [57, 58, 59]. Their synthesis also tends to yield more complex and less uniform shapes, adding to the spectral diversity. Despite the intrinsically higher variability of NCoM spectra (SI Fig. S6(b)) compared to NPoMs (Fig. 3(d)), SPARX consistently achieved substantially lower prediction errors than a naive statistical baseline (SI Fig. S6(c)), demonstrating that SPARX serves as a general framework capable of handling diverse plasmonic nanostructures beyond spherical geometries.
@@
Discussion
A key bottleneck in Figure 5 is the exposure time of our RGB camera (2s per snapshot for high resolution), as SPARX’s inference itself takes only fractions of a second (Figure 5(b)). Future improvements in throughput and performance can be approached from both algorithmic and optical perspectives. Algorithmically, SPARX inference is already fast, but accuracy and uncertainty estimation could be enhanced by refining the heteroscedastic loss function, exploring alternative error distributions (e.g., heavy-tailed or skewed), optimizing network architectures and hyperparameters, or adopting advanced models such as variational autoencoders (VAEs) or Bayesian neural networks (BNNs) to better capture stochastic single-particle spectral responses. Optically, acquisition speed could be improved by increasing camera gain (leveraging the model’s noise robustness), expanding the field of view with lower-magnification objectives or wide-field illumination, boosting illumination intensity using high-brightness sources, and enhancing spatial resolution with higher NA objectives, confocal configurations, or advanced techniques such as structured illumination microscopy [60], super-resolution plasmonics [61], or DF nanoscopy [62]. These combined algorithmic and optical strategies can significantly accelerate measurements while maintaining or improving spectral prediction accuracy.
From the perspective of information acquisition, spectral optimization presents two complementary avenues. Firstly, one can tailor the plasmonic systems to exhibit dominant resonant peaks within the visible range (400–700 nm). It would align spectral signatures with the RGB camera’s detection window, maximizing signal capture efficiency and allowing significantly shorter exposure times. Secondly, inspired by hyperspectral techniques [63], a few detection channels beyond 700 nm can be incorporated—for example, via multi-channel split-frequency systems such as grayscale CCDs with spectral filters—to broaden the spectral window up to 1000 nm or beyond, thereby capturing near-infrared resonances.
@@ Discussion
Beyond spectral reconstruction, DF RGB images encode additional information that SPARX can exploit for nanoparticle shape classification. Since metallic nanoparticle’s shape strongly influences plasmonic performance [64], and synthesis often yields by-products with varied geometries, this approach provides a valuable, noninvasive alternative to conventional techniques such as scanning electron microscopy (SEM), transmission electron microscopy (TEM), and atomic force microscopy (AFM). In future applications, SPARX classifier could distinguish multiple particle geometries simultaneously on the same substrate, enabling additional rapid screening of desired nanoparticles or studies that leverage spectral diversity from different particle types [65].

### v3

@@ Abstract
keywords: Deep learning, Dark-field, Near-field, Far-field, Single plasmonic nanocavity, Consistency, Predicting spectra from images, Classification
@@ Introduction
Plasmonic nanostructures, particularly those forming nanogaps, support localized surface plasmons that confine light to deep-subwavelength volumes comparable to molecular and atomic dimensions. 63, 21, 5, 7, 14. By integrating low-dimensional materials into these nanogaps, they enable probing of electronic (e.g., photoluminescence) 2, vibrational (e.g., surface-enhanced Raman scattering) 34, 7, and other linear and nonlinear optical properties of materials 10, 11, with atom-level sensitivity 18, introducing a fascinating field as “extreme nanophotonics5, 40. This high sensitivity arises from the extremely enhanced nanoscopic light-matter interactions supported by nanostructure’s quasinormal modes 39, making resonance–material spectral matching a fundamental prerequisite for nanophotonic applications.
@@
However, this extreme localization also amplifies contributions from local nanomorphological features such as nonuniformity, roughness, defects, surface atomic-level dynamics, and pico-cavities inevitably introducing ‘noise’ and compromising reproducibility 59, 15, 62, 56, 39, 19, 7, 6, 57. This intrinsic variability 26, 5 poses a paradox—enhancing fields for exceptional single-particle performance can inherently undermine spectral consistency and reproducibility across ensembles, constituting a crucial bottleneck in extreme nanophotonics. One general way of circumventing this is to identify and pre-screen the target (e.g., spectrally-uniform, with desired resonances) nanoparticles from countless particles ($\simeq 10^{10}$ per milliliter in the solution) drop-casted on the substrate during the sample preparation, and conduct experiments only on that subset. Conventionally, this screening process has relied on manual expertise combined with imaging and spectroscopic measurements, rendering it inefficient and labor-intensive.
@@
Confocal dark-field (DF) microscopy has long been a cornerstone for single-nanoparticle characterization. Resonances and sample uniformity are often inferred empirically from their images at high throughput, assuming that the color distribution of Airy patterns directly reflects the energy and lineshape of the plasmonic modes. However, human vision cannot reliably resolve subtle chromatic variations, and the information-compressed three-digit RGB output of commercial cameras (400–700 nm) lacks sufficient spectral detail and fails to capture lower-energy resonances. Multiple resonances with distinct far-field patterns further complicate the problem 51, 33. More precise techniques, such as DF mapping, acquire spatially resolved spectra via point-scanning, representing a form of hyperspectral imaging (HSI) 16, 67, 48, 17, 43, 8. Such sequential acquisition is unsuitable for large-field detection, which can be mitigated with alternative HSI strategies, including line-scanning (push-broom), wavelength-scanning 16, and snapshot techniques 8. Recent advances, including compressive sensing 50, 64, 23, 65, have further improved acquisition speed. However, HSI methods inherently face a resolution–throughput trade-off: higher spectral resolution requires narrower channels, reducing photon flux per channel 58, 43. This lowers the signal-to-noise ratio and, unless compensated by stronger illumination or more sensitive detectors, demands longer exposure times which is especially critical for weak signals such as single nanoplasmonic structures. They also entail costly and specialized hardware, large data storage and copious processing requirements, and rigorous calibration protocols 67, 1. In contrast, confocal DF microscopy combined with RGB imaging remains economical and effective, while readily compatible with emerging techniques (e.g., machine learning, compressive sensing). However, to fully leverage this approach for high-throughput, precise acquisition of plasmonic nanostructures, the complex and nonlinear correlations between DF images and spectra, particularly beyond the RGB camera’s physical limits, must be revealed.
@@
In this regard, deep learning (DL) has emerged as a transformative tool for decoding complex photonic interactions 61. It has demonstrated the ability to learn intricate light–matter correlations through inverse design of plasmonic devices and metasurfaces 44, 28, 42, 41, 3, 53, 4, 31. When integrated with microscopic and nanoscopic techniques, DL has also shown success in noise suppression, signal processing, nanoparticle identification, and super-resolution imaging, which resolves variations and uncertainties at the subwavelength scale 66, 54, 25, 22, 27, 29, 30, 55, 38, 37. Overall, DL is particularly valuable for managing uncertainties in experimental and fabrication processes, such as material inconsistencies, structural imperfections, and environmental fluctuations.
@@
Furthermore, by leveraging heteroscedastic loss, the SPARX model could estimate the uncertainty in spectral predictions for individual nanoparticles, effectively quantifying the prediction error. This correlation allows for further refinement in the selection process. Moreover, as a powerful complementary screening approach, we develop a SPARX classifier capable of directly identifying and differentiating nanoparticle shapes from DF RGB images, showcasing an inference capability beyond the reach of conventional spectral acquisition methods. All in all, our approach achieves a $10^{2}$$10^{4}$-fold acceleration in characterization speed compared with spectral acquisition, enabling real-time classification of nanostructures in the future. By replacing spectrometer-dependent workflows with camera-based deep learning, we establish a scalable platform for next-generation optical characterization, nanophotonic device engineering, and biosensor development, where both spectral precision and high-throughput characterization have proven indispensable.
Results and Discussion
@@
Among all plasmonic systems, the nanoparticle-on-mirror (NPoM) configuration (Figure 1(a)) has been widely observed as one of the most versatile platforms, offering easily controllable nanogaps for enhancing optical phenomena 5. Therefore, we select NPoM as a representative platform for demonstration, measuring 12,000 gold NPoMs to compile a dataset that captures the underlying complexity. In our experiment, the NPoM system was formed by drop-casting 80 nm cetyltrimethylammonium (CTAC)-capped gold nanoparticles onto a gold mirror (see Methods for sample preparation). The nanogap was defined by a 1-2 nm thick CTAC molecule layer. From electromagnetic theory, it is well established that the size and shape of the nanoparticles 62, 6, small variations in the roughness of the metallic surface 15, 59, and the morphology of the gap between the nanoparticles and the metallic substrate (e.g., facets size and adatoms) 57, 39 can drastically alter the resonance behavior of the system.
To visualize these effects, we performed finite element method (FEM) simulations for the DF scattering spectra of the NPoMs using the commercial software COMSOL Multiphysics (see Methods for details). To avoid exhaustive geometric cases discussed above, our simulations focus on two primary geometric factors varying at the nanometer scale: the facet size, and the nanogap thickness between the nanoparticle and the substrate. The results, presented in Figure 1(b), clearly show that even minor modifications to these geometrical parameters lead to significant shifts in the resonances. Increasing the facet size may redshift the resonances, which can be related to the polyhedral morphology of the metal’s crystalline structure and light-induced atomic migration 62. The nanogap thickness is another crucial factor that may substantially shift the resonances, since the nanogap plasmons have been proven to be able to resolve sub-picometer level of thickness variation 12. In our test sample, the gap thickness can inherently vary on the level of a few angstroms.
@@
The attempt to uncover the correlations among different orders of the plasmonic resonances can be dated back to the Mie theory 9, where an analytical solution of the scattering from a spherical nanoparticle can be derived by solving the Maxwell equations. Each Mie resonance (i.e., dipole, quadrupole, etc.) can be intercorrelated through the spherical harmonics. When coupled with its own image through a mirror, the NPoM plasmonic nanogap can give rise to a richer variety of the hybridized modes 52, with the resonance of $mn$-order mode $\lambda_{mn}$ given by the cylindrical Bessel functions 57: $\lambda_{mn}=\frac{\pi wn_{\mathrm{eff}}}{J_{mn}-\phi}$, where $J_{mn}$ is $n$-th root of the $m$-th order Bessel function. $n_{\rm eff}$ is the effective refractive index of the metal-insulator-metal junction, and $\phi$ is a proper reflection phase. However, this simplified model can only predict the cavity-like resonances of a perfectly spherical nanoparticle with a single ideal circular facet at the bottom gap, with uncertainties such as polyhedral shapes, roughness, nonlocal and quantum effects, and others mentioned above falling outside its scope. Thus, using a simplified, classical analytical solution to extend the spectra of a realistic system beyond the detection limit can be challenging.
In contrast, although certain dark-field images may appear visually similar, they remain distinguishable in the high-dimensional feature space used by the model. Quantitative comparisons of visually similar image pairs and their corresponding spectra are presented in Fig. S10 in the Supplementary Information. As given sufficiently representative datasets, DL methods excel at resolving complex and nonlinear correlations with high degrees of freedom. In fact, our SPARX model has managed to decode the broadband spectrum varying from 500-1000 nm from the images and therefore could select the spectra with the minimal predicted MAE compared with the reference. It hit 3 out of 10 actual best choices as shown in Figure 1 (c), breaking away from the color-based selection rule. Nonetheless, the rest 7 “missed” selections also have a great match with the reference peak, showing a nearly overlapped mean value (Figure 1 (d)) with low standard deviation. Additionally, Figure 1(e) summarizes the mean, standard deviation, minimum, and maximum values of the resonance peak location differences with respect to the reference particle. Again, this analysis shows that the SPARX’s selection closely aligns with the actual best selection, exhibiting similar shifts and variances. This reinforces the capability of deep learning in reliably identifying and predicting nanoparticles with optimal spectral properties, outperforming human intuition. Next, we will explain the process of preparing the datasets and training the SPARX model.
@@
We perform principal component analysis (PCA) on both the DF images and the processed spectra. PCA reduces high-dimensional data by projecting it onto new orthogonal axes, principal components, that capture most of the variance (i.e., the directions in which the data varies the most or has the most spread). See Methods for data processing and multivariate analysis for more details. The variance explained by the first 128 PCA components for both datasets is shown in Figure 2(b). Figures 2(c) and (d) display the first four PCA components of the DF images and spectral data, respectively. In Figure 2 (c), each PCA component is presented with its corresponding red, green, and blue channels. A notable observation is the correlation between the size of extracted features and the diffraction limit of each color channel. For instance, in PC1 of the DF image (Figure 2(c), top row), the dark central region in RGB combined plot (where the data is near zero) appears in different colors across the RGB image, but its size decreases from red to blue, consistent with diffraction-limited resolution. Similar trends are observed in other components, such as the concentric ring structures in PC2, which also exhibit size reduction across color channels. On the other hand, the patterns of the PC1–4 for the DF images clearly present the spherical harmonics (e.g., $l=0$ for PC1, $l=1$ for PC2–4), inferring the NPoM’s point scattering nature in the Airy pattern 51. These observations demonstrate that our analysis based on PCA captures meaningful physical features, revealing both the optical and structural characteristics of the system. Furthermore, using the PCA of DF images we are able to access scattering information and potentially spectral signatures.
@@
From the spectral PCA components in Figure 2(d), PC1 which captures the largest amount of variance primarily reflects peak shift behavior. It shows higher values when the spectral peak shifts toward the red (1000 nm) or blue (400 nm) regions, with a local minimum around 800 nm. This behavior aligns with the simulations in Figure 2(b), where minor geometric variations, such as changes in facet size or nanogap thickness, lead to significant spectral shifts. The remaining PCs capture finer spectral details, with alternating positive peaks and negative dips across various wavelengths. Thus, our analysis based on the PCA of DF spectra not only reveals spectral variations but also encodes geometric information.
@@
To explore the correlation between DF images and spectra, we project both datasets into lower-dimensional PCA spaces and visualize the results in Figure 2(e). According to Mie and plasmon hybridization theories, the visible higher-order modes and the near-infrared gap-plasmon modes in NPoM systems originate from the same hybridized eigenmode family, so variations in gap thickness or facet size mainly lead to correlated spectral shifts rather than mode decoupling. This correlation can be evidenced by the data-driven multivariate analysis. The density distribution of the spectral data along PC1 reveals multiple subpopulations, suggesting that this component captures distinct spectral characteristics. In the 2D PCA projection of DF images, two well-separated clusters emerge (middle panel of Figure 2 (e)). When coloring the DF image projections using different criteria (see Methods: Multivariate Analysis), clear patterns can emerge. First, using spectral PC1 projection reveals an alignment between DF image features and spectral variation. Second, using sample indices, reflecting the acquisition timeline, demonstrates the consistency of measurements over time. Third, spectral features, namely the peak location (wavelength $\lambda$) and total spectral energy, further validate that the image-based clustering corresponds to meaningful spectral characteristics. These patterns highlight the physical relevance of the clusters found in our analysis (Figure 2(e))
@@
To further support these observations, we applied Uniform Manifold Approximation and Projection (UMAP), a nonlinear dimensionality reduction technique, to both spectral and DF image data, as shown in SI Figure S2. The spectral data is embedded into a one-dimensional UMAP space, while the DF images are projected into two dimensions. By coloring the DF UMAP projection using various spectral-derived features including UMAP values of the spectra, peak positions, and total spectral energy, we uncover strong visual correlations between the DF and spectral modalities (2D maps in SI Figure S2). An interesting outcome is the appearance of multiple distinct clusters in the DF UMAP space (in contrast to PCA projection in Figure 2 (e)), which exhibit clear correspondence to specific spectral features. For example, the cluster located at the bottom right of the projection space shows a color gradient where resonance wavelength and total spectral energy vary in opposite directions, i.e. as the resonance red-shifts, the intensity decreases. Additionally, when coloring the projections based on sample index, correlated with the data acquisition time, we observe no discernible pattern, confirming the high reproducibility and stability of the measurements across multiple days. This improved clustering allows us to reveal the rich spectral information that DF images inherently carry, which is typically inaccessible in traditional DF imaging analysis.
@@
In contrast, the intensity distribution, shown in the right panel of Figure 3(b), presents a distinctly different behavior. The overall population density plot of resonance intensity (shown in black) displays three distinct peaks, indicating the existence of three subpopulations. Interestingly, the core group is primarily composed of the subpopulation with the lowest resonance intensities. The remaining two subpopulations with higher intensities are predominantly classified as outliers. Interestingly, the spectra with modest intensity ($\simeq$2.6 arb. units), which contain the largest number of samples, are most recognized as the core data that SPARX is most confident in. This may be attributed to their generally more uniform and well-behaved geometries, with reduced randomness stemming from previously discussed factors. In contrast, the higher-intensity subpopulations are treated as outliers by SPARX, likely due to more complex features such as sharp polyhedral facets, nonuniform gaps, or even picocavities. These features have more randomness in general, and are more difficult to characterize using only DF images, leading to increased uncertainty in their prediction. That means, the identification of outliers is primarily data-driven rather than based on direct structural evidence from SEM or TEM, since the features inside the nanogap are typically inaccessible by SEM or TEM. In fact, this also aligns with the empirical intuition in plasmonics and surface-enhanced Raman scattering: extreme enhancement with high intensity often compromises robustness and consistency. Nanoparticles exhibiting such exceptionally high intensities remain rare outliers relative to the overall population. Remarkably, our SPARX model captures this trade-off by analyzing the associated uncertainty. It is important to note that while our heteroscedastic model is capable of identifying such complex cases from DF images, accurately reconstructing their spectral details remains challenging due to the limitations of the input modality. Detection of these outliers should not be conflated with precise spectral prediction.
@@
Beyond NPoM, we further validated the generality of SPARX on a more complex plasmonic system that is nanopatch antennas. This system is composed of single $\sim$75 nm gold nanocubes on a mirror (i.e., NCoM, see SI Section S5, Figs. S6, S7). Unlike nanospheres, nanocubes produce more diverse spectral responses, as geometric variations can generate multiple resonances associated with different gap sizes and bottom facet geometries 35, 49, 13. Their synthesis also tends to yield more complex and less uniform shapes, adding to the spectral diversity. Despite the intrinsically higher variability of NCoM spectra (SI Fig. S6(b)) compared to NPoMs (Fig. 3(d)), SPARX consistently achieved substantially lower prediction errors than a naive statistical baseline (SI Fig. S6(c)), demonstrating that SPARX serves as a general framework capable of handling diverse plasmonic nanostructures beyond spherical geometries.
@@
A key bottleneck in Figure 5 is the exposure time of our RGB camera (2s per snapshot for high resolution), as SPARX’s inference itself takes only fractions of a second (Figure 5(b)). Future improvements in throughput and performance can be approached from both algorithmic and optical perspectives. Algorithmically, SPARX inference is already fast, but accuracy and uncertainty estimation could be enhanced by refining the heteroscedastic loss function, exploring alternative error distributions (e.g., heavy-tailed or skewed), optimizing network architectures and hyperparameters, or adopting advanced models such as variational autoencoders (VAEs) or Bayesian neural networks (BNNs) to better capture stochastic single-particle spectral responses. Optically, acquisition speed could be improved by increasing camera gain (leveraging the model’s noise robustness), expanding the field of view with lower-magnification objectives or wide-field illumination, boosting illumination intensity using high-brightness sources, and enhancing spatial resolution with higher NA objectives, confocal configurations, or advanced techniques such as structured illumination microscopy 24, super-resolution plasmonics 60, or DF nanoscopy 36. These combined algorithmic and optical strategies can significantly accelerate measurements while maintaining or improving spectral prediction accuracy.
From the perspective of information acquisition, spectral optimization presents two complementary avenues. Firstly, one can tailor the plasmonic systems to exhibit dominant resonant peaks within the visible range (400–700 nm). It would align spectral signatures with the RGB camera’s detection window, maximizing signal capture efficiency and allowing significantly shorter exposure times. Secondly, inspired by hyperspectral techniques 68, a few detection channels beyond 700 nm can be incorporated—for example, via multi-channel split-frequency systems such as grayscale CCDs with spectral filters—to broaden the spectral window up to 1000 nm or beyond, thereby capturing near-infrared resonances.
@@
Beyond spectral reconstruction, DF RGB images encode additional information that SPARX can exploit for nanoparticle shape classification. Since metallic nanoparticle’s shape strongly influences plasmonic performance 46, and synthesis often yields by-products with varied geometries, this approach provides a valuable, noninvasive alternative to conventional techniques such as scanning electron microscopy (SEM), transmission electron microscopy (TEM), and atomic force microscopy (AFM). In future applications, SPARX classifier could distinguish multiple particle geometries simultaneously on the same substrate, enabling additional rapid screening of desired nanoparticles or studies that leverage spectral diversity from different particle types 20.